# Towards a Cascaded LLM Framework for Cost-effective Human-AI Decision-Making

### v2

@@ Abstract
Effective human-AI decision-making balances three key factors: the correctness of predictions, the cost of knowledge and reasoning complexity, and the confidence about whether to abstain automated answers or involve human experts. In this work, we present a cascaded LLM decision framework that adaptively delegates tasks across multiple tiers of expertise – a base model for initial candidate answers, a more capable and knowledgeable (but costlier) large model, and a human expert for when the model cascade abstains. Our method proceeds in two stages. First, a deferral policy determines whether to accept the base model’s answer or regenerate it with the large model based on the confidence score. Second, an abstention policy decides whether the cascade model response is sufficiently certain or requires human intervention. Moreover, we incorporate an online learning mechanism in the framework that can leverage human feedback to improve decision quality over time. We demonstrate this approach to general question-answering (ARC-Easy and ARC-Challenge) and medical question-answering (MedQA and MedMCQA). Our results show that our cascaded strategy outperforms in most cases single-model baselines in accuracy while reducing cost and providing a principled way to handle abstentions. [^1] We provide the code for our experiments at https://github.com/fanconic/cascaded-llms
Machine Learning, ICML
@@ 1 Introduction
Data-driven decision support has gained increasing traction in high-stakes fields such as healthcare (Jin et al., 2024; Fan et al., 2024; Li et al., 2024), finance (Li et al., 2023a; Zhao et al., 2024), and education (Xu et al., 2024). For example, in the medical context, large language models (LLMs) can facilitate accurate diagnoses and treatment recommendations that encode vast knowledge (Kim et al., 2024). However, high accuracy in such complex settings often requires substantial computational resources or multiple reasoning steps. Additionally, LLMs may hallucinate or generate incorrect outputs with severe consequences. Effective human-AI collaboration should balance correctness, cost, and abstention, ensuring AI-driven assistance integrates seamlessly with expert oversight.
Challenge:
Efficient allocation of computational and human resources is essential, as larger models incur significantly higher costs, especially with advanced reasoning methods, like searching the solution space (OpenAI, 2025; Anthropic, 2025). Decision-making tasks differ in complexity. Smaller models can handle routine cases, while complex queries require more advanced reasoning. Choosing between deploying a small model, a more sophisticated model, or referring to a human expert is non-trivial, particularly under resource constraints. A static approach — always using a large model or defaulting to small ones — fails to optimise cost-benefit trade-offs.
@@
Figure 1: Cascaded LLM Human-AI Decision-Making Framework. Given a decision-making problem, the system (1) generates an initial response with a base model, (2) verifies correctness probability, (2.5) defers to a larger model if needed, (3) assesses response uncertainty, and (3.5) abstains to a human expert if necessary. If feedback is available, deferral and abstention modules are adjusted over time. For this system to work efficiently, the modules should uphold three desiderata: D1 the deferral policy regenerates responses only when necessary, D2 the abstention policy escalates to humans only when uncertainty is high, D3 the system continuously improves with feedback.
Our Solution:
We propose a cascaded multi-tier framework integrating (1) a cost-effective base model, (2) an advanced but resource-intensive large model, and (3) a human expert in case of abstention. This system prioritises efficiency, deferring only when likely and necessary. A deferral policy determines if the large model is required to regenerate the answer, while an abstention policy forwards the query to experts when uncertainty is high. Online learning continuously refines these policies based on feedback. Figure 1 illustrates how decisions are handled across datasets of varying difficulty.
Framework Desiderata:
Three desiderata must be upheld for our proposed framework to function correctly:
D1 : Reduce Unnecessary Regenerations The deferral policy should escalate to the large model only if the base model’s response is likely incorrect, avoiding unnecessary computational costs.
D2 : Abstain and Escalate to Humans when Uncertain The abstention policy should refer to human experts only when the cascade system’s uncertainty is too high, reducing unnecessary human intervention.
D3 : Dynamic Improvement based on Feedback The framework should continually refine its policies based on feedback, ensuring adaptive and risk-minimised decision-making.
Contributions:
We make the following key contributions in this paper:
  • Cascaded LLM human-AI decision system: A framework integrating LLMs of varying complexity with human experts for accuracy, efficiency, and uncertainty management.
  • Investigation into Cost-Efficient Verification Methods: We analyse different LLM verification methods to determine the most suited for cost-efficient deferral.
  • Online learning framework: An online learning approach that continuously improves deferral and abstention policies when feedback is available, enhancing decision-making adaptively over time.
@@ 2 Related Work
Multi-LLM Answer Generation. Several studies have explored collaborative frameworks that leverage multiple LLMs of varying capacities to enhance both performance and cost-efficiency beyond the capabilities of a single model (Chen et al., 2023; Ding et al., 2024; Aggarwal et al., 2024). (Chen et al., 2023) proposed cost-effective strategies such as prompt structuring, model approximation, and cascaded LLM frameworks. Similarly, (Ding et al., 2024) introduced an intelligent routing mechanism that dynamically assigns prompts to the most appropriate model. (Aggarwal et al., 2024) developed a black-box LLM framework for cost-efficient response generation, formalised as a Partially Observable Markov Decision Process (POMDP), requiring minimal training data. (Zhu et al., 2023a) proposed a multiplexer-based approach that balances queries between a small and a large LLM, employing a trained BERT classifier to determine when the smaller model suffices. (Šakota et al., 2024) introduced a meta-model-driven selection framework that requires pre-training for optimal query distribution. In a parallel line of research, speculative decoding (Leviathan et al., 2023), employs a lightweight model to generate multiple tokens, which a larger model subsequently verifies.
@@
In contrast to prior research, we propose a multi-tier framework for human-AI collaboration. Rather than relying solely on automation, our approach integrates human intervention when model uncertainty is too high, addressing a gap in previous multi-tier frameworks. Compared to speculative decoding research, our work prioritises the factual correctness of complete responses rather than token-wise distributions, enabling more robust decision-making rather than just fluent text generation. (Zellinger et al., 2025) conducts a concurrent line of research that is closest to our work on cascaded LLMs, as well as in their previous works (Zellinger and Thomson, 2024, 2025). They focus on probabilistic modelling of cascading LLMs and their deferral and abstention mechanisms.
@@
LLM Answer Verification and Uncertainty Quantification. Ensuring the reliability of LLM-generated responses requires adequate verification and uncertainty quantification mechanisms. Several studies have explored self-verification strategies (Weng et al., 2023; Jiang et al., 2024; Pan et al., 2024), often leveraging the LLM’s internal knowledge (Dhuliawala et al., 2023). Alternative approaches employ external knowledge sources for verification (Pan et al., 2024; Gao et al., 2023; Peng et al., 2023). (Aggarwal et al., 2024) introduced verification techniques based on available contextual information, predominantly involving multiple LLM queries to validate response accuracy. Another research direction quantifies factual correctness uncertainty (Mahaut et al., 2024). (Kadavath et al., 2022) conducted a detailed analysis of how LLMs express uncertainty through surrogate token probabilities, demonstrating their effectiveness in calibration. (Azaria and Mitchell, 2023) explored internal LLM states, training classifiers to quantify uncertainty, while methods such as semantic uncertainty estimation (Kuhn et al., 2023) enhance robustness by analysing variations in semantically equivalent token sequences.
@@
LLMs in Online Learning. Traditional LLM research predominantly evaluates language models on static datasets. However, our work aligns with online learning paradigms, wherein policies are continuously refined in response to streaming data (Cortes et al., 2018; Ye et al., 2024). Our methodology is inspired by (Jarrett et al., 2022), who introduced an online decision mediation framework mediating between suboptimal human decisions and an expert oracle. A similar research with the online learning approach is conducted by (Zhu et al., 2023a), which extended their multiplexer mechanism to an online setting.
@@ 3.1 Cascaded Decision System
$\tiny C(x)=\begin{cases}\mathcal{M}_{\text{base}}(x)&\text{if }\Phi_{\text{base}}(x)>\phi_{\text{base}}\land\Xi_{\text{base}}(x)<\xi_{\text{base}}\\ \mathcal{M}_{\text{large}}(x)&\text{if }\Phi_{\text{base}}(x)\leq\phi_{\text{base}}\land\Xi_{\text{base}}(x)\leq\xi_{\text{base}}\land\Xi_{\text{large}}(x)\leq\xi_{\text{large}}\\ \varnothing&\text{if }\Xi_{\text{base}}(x)\geq\xi_{\text{base}}\,\lor\,\Xi_{\text{base}}(x)\geq\xi_{\text{base}}\end{cases}$ (1)
@@
Figure 2: Decision flow of the two-tiered cascaded LLM system. A query is first evaluated by the base model. The base model response is returned if the confidence is above a deferral threshold and uncertainty is below an abstention threshold. Otherwise, the query is passed to the large model, producing a response or abstaining if uncertainty remains high. In such cases, the query is deferred to a human expert. Online feedback can be incorporated at each stage to improve system performance over time.
@@ 3.2 Cost Calculation
To estimate the computational cost incurred by a language model when generating a response, we define a cost function that scales linearly with the number of parameters and the number of tokens processed. Let $s\in\mathbb{R}_{>0}$ denote the size of the model in billions of parameters, extracted from the model identifier (e.g., “Llama-8B” $\Rightarrow$ $s=8$). Let $t_{\text{in}}$ and $t_{\text{out}}$ denote the number of input and output tokens, respectively, and let $\rho>0$ denote the output-to-input token price ratio, which accounts for the higher cost of generating tokens. The total cost is given by:
@@
This abstraction provides a simple yet effective means of comparing models of different sizes and outputs under a unified cost metric, without relying on specific infrastructure-dependent pricing.
@@ 4.1 Calibrated Confidence and Uncertainty Estimation
For each of the above methods, the evaluating model $\mathcal{M}_{i}$ can either be the same model that generated the original response, or a larger model in the cascade, if available. While self-evaluation is cheap and self-contained, verifying a small model’s output using a larger model is still substantially cheaper than generating a new response from scratch—particularly when generation involves long-form reasoning, as per Assumption  A2. Additionally, larger models tend to be better calibrated and may yield more reliable verification, improving downstream deferral and abstention decisions (Zhu et al., 2023b; Chhikara, 2025).
@@
To ensure that the extracted confidence scores are comparable across models and consistent with empirical correctness, we fit a Bayesian logistic regression model on a small calibration set of 100 samples. This is a Bayesian version of Platt scaling (Platt, 2000), and we assume a Normal distribution as prior. We follow (Zellinger and Thomson, 2024)’s approach and apply a non-linear transformation on the raw confidence score before inputting it into the Bayesian model, to spread out the clusters of overconfident probabilities.
@@ 4.2 Online Improvement
$\mathbb{P}(\text{abstention})\;=\;p_{\text{abst1}}\;+{p_{\text{abst2}}}.$
@@
$\mathbb{E}[\text{Correct}]\;=\;\mathbb{E}\!\bigl{[}m_{\text{pred1}}\cdot\Phi_{\text{base}}\bigr{]}\;+\;\mathbb{E}\!\bigl{[}m_{\text{pred2}}\cdot\Phi_{\text{large}}\bigr{]}.$
@@
$\mathbb{E}[\text{Cost}]\;=\;c_{1}\;+\;\mathbb{E}[m_{\text{defer1}}]\cdot c_{2}.$
@@
(a) ARC2-Easy
(b) ARC2-Challenge
(c) MedQA
(d) MedMCQA
Figure 3: Cost-Accuracy Trade-off for Calibrated Verification Methods (Llama 1B–8B). Accuracy versus cost per sample is shown for the cascaded model using various verification methods. Performance above the linear interpolation line between base and large model baselines indicates a positive cost-benefit. Error bars represent standard error.
@@ 5 Experiments
To evaluate the generalisability of our framework across domains, we use four question-answering datasets: (1) ARC2-Easy and (2) ARC2-Challenge (Clark et al., 2018), which are part of the AI2 Reasoning Challenge and require reasoning over grade-school science; and two medical QA benchmarks: (3) MedQA (Jin et al., 2020), consisting of US medical board exam questions, and (4) MedMCQA (Pal et al., 2022), comprising entrance exam questions from the Indian medical school curriculum. All datasets are in multiple-choice format, with ground-truth answers satisfying Assumption A3. Chain-of-Thought reasoning is employed to generate answers. The cost proportion between input and output tokens is set to $\rho=5$, consistent with Anthropic’s current pricing to date (Anthropic, 2025). Details on generation and verification prompts can be found in Appendix B.1. The codebase is available at https://github.com/fanconic/cascaded-llms.
@@ 5.1 Cost-Benefit Analysis of Verification Methods
To assess both cost-efficiency and accuracy, we compare the calibrated base model confidence $\Phi_{\text{base}}$ against two baselines: (1) using only the base model (Llama-1B) and (2) using only the large model (Llama-8B). In Figure 3, we visualise accuracy versus cost per sample across the datasets. We use a threshold-agnostic strategy where deferral to the large model is performed with probability $\Phi_{\text{base}}(x)$. We evaluate four methods: Self-Verification (SV, $n{=}1$), Surrogate Token Probability (STP, $n{=}1$), Consistent Self-Verification (SV, $n{=}5$), and Monte Carlo STP (MC-STP, $n{=}5$). For the latter two, we perform five regenerations or stochastic passes. Each experiment is conducted once using $\mathcal{M}_{\text{base}}$ as the verifying LLM, and once using $\mathcal{M}_{\text{large}}$.
As shown in Figure 3 (and Figures 6, 9, and 12 in Appendix C), using a larger model for verification generally yields a better cost-benefit profile, particularly on simpler datasets (ARC2-Easy, ARC2-Challenge). In contrast, base-model verification provides only marginal gains. On the harder medical datasets (MedQA and MedMCQA), all methods struggle. STP ($n{=}1$) is the most effective in subfigures 3(a) and 3(b).
ARC2-Easy ARC2-Challenge MedQA MedMCQA
Base SV ($n{=}1$) -8.8 $\pm$ 16.4 -13.2 $\pm$ 11.4 -8.0 $\pm$ 9.5 -9.9 $\pm$ 8.9
SV ($n{=}5$) -12.8 $\pm$ 16.5 -19.6 $\pm$ 10.8 -21.3 $\pm$ 8.6 -15.6 $\pm$ 8.7
STP ($n{=}1$) -4.4 $\pm$ 16.7 -5.2 $\pm$ 11.7 -13.6 $\pm$ 9.3 -4.8 $\pm$ 9.3
STP ($n{=}5$) 0.8 $\pm$ 15.2 -9.0 $\pm$ 11.4 -9.4 $\pm$ 9.2 -11.2 $\pm$ 8.8
@@
SV ($n{=}5$) 50.1 $\pm$ 13.7 10.7 $\pm$ 11.8 -8.7 $\pm$ 9.1 -3.1 $\pm$ 9.0
STP ($n{=}1$) 118.3 $\pm$ 15.8 38.1 $\pm$ 13.7 5.6 $\pm$ 9.7 -3.7 $\pm$ 9.2
STP ($n{=}5$) 97.7 $\pm$ 14.4 15.4 $\pm$ 11.7 -6.8 $\pm$ 9.2 2.6 $\pm$ 9.0
Table 1: $\Delta$IBC Scores for Calibrated Cascade (Llama 1B–8B). Each row indicates a verification method (SV or STP) with $n=1$ or $n=5$, grouped by whether the base or large model was used for verification. A higher score indicates better cost-efficiency relative to the baseline.
@@
As seen in Table 1 verifying with $\mathcal{M}_{\text{large}}$ consistently leads to higher $\Delta$IBC scores, particularly on ARC2-Easy and ARC2-Challenge. On the medical datasets, no single method consistently outperforms the others significantly. Moreover, we see that on the medical datasets, the $\Delta$IBC standard error rates for the verification scores using the large are around 0, indicating no cost-benefit compared to the easier datasets. We report additional results for the other cascades in Tables 2,  3,  4, and 5 in Appendix C. Interestingly, the uncalibrated confidence scores seem to lead to higher $\Delta$IBC, but with a much higher standard error, indicating the instability of uncalibrated confidence scores.

### v3

# Cascaded Language Models for Cost-Effective Human–AI Decision-Making

@@ Abstract
A challenge in human-AI decision-making is to balance three factors: the correctness of predictions, the cost of knowledge and reasoning complexity, and the confidence about whether to abstain from automated answers or escalate to human experts. In this work, we present a cascaded LLM decision framework that adaptively delegates tasks across multiple tiers of expertise – a base model for initial candidate answers, a more capable and knowledgeable (but costlier) large model, and a human expert for when the model cascade abstains. Our method proceeds in two stages. First, a deferral policy determines whether to accept the base model’s answer or regenerate it with the large model based on the confidence score. Second, an abstention policy decides whether the cascade model response is sufficiently certain or requires human intervention. Moreover, to overcome static policies and accommodate changing task difficulty, we incorporate an online learning mechanism which uses human feedback. We demonstrate this approach to general question-answering (ARC-Easy, ARC-Challenge, and MMLU) and medical question-answering (MedQA and MedMCQA). Our results demonstrate that our cascaded strategy outperforms single-model baselines in most cases, achieving higher accuracy while reducing costs and providing a principled approach to handling abstentions. [^1] We provide the code for our experiments at https://github.com/fanconic/cascaded-llms
@@ 1 Introduction
Data-driven decision support has gained increasing traction in high-stakes fields such as healthcare (Jin et al., 2024; Fan et al., 2024; Li et al., 2024), finance (Li et al., 2023a; Zhao et al., 2024), and education (Xu et al., 2024). For example, in the medical context, large language models (LLMs) can facilitate accurate diagnoses and treatment recommendations that encode vast knowledge Kim et al. (2024). However, high accuracy in such complex settings often requires substantial computational resources or multiple reasoning steps. Additionally, LLMs may hallucinate or generate incorrect outputs with severe consequences. Effective human-AI collaboration should balance correctness, cost, and abstention, ensuring AI-driven assistance integrates seamlessly with expert oversight.
The Challenge.
A key challenge in effective human–AI collaboration is how to allocate computational and human resources efficiently—deciding when an automated model should answer, when it should escalate to a lager and more capable model, and when it should defer to human expertise. Naïve strategies — such as always relying on the cheaper model or always trusting the more capable one — fail to optimise this trade-off. The former increases the risk of errors and hallucinations, while the latter inflates costs. Likewise, static deferral policies, fixed thresholds, or one-off calibrations cannot adapt to changing task distributions or evolving model competence.
These limitations motivate three requirements for any cost-effective human–AI decision-making framework:
  1. 1.
    D1
    Reduce Unnecessary Regenerations: Responses should only be regenerated by a more capable model when there is sufficient evidence that the current one is unreliable..
  2. 2.
    D2
    Abstain when Uncertain: The system should defer to human experts when uncertainty exceeds acceptable bounds, avoiding overconfident automation in high-risk scenarios.
  3. 3.
    D3
    Adapt over time: The framework should continuously refine its deferral and abstention policies as feedback becomes available, ensuring sustained reliability and improvement.
Together, these desiderata define the principles an effective decision-making framework must satisfy, irrespective of implementation.
Our Approach.
We propose a cascaded LLM framework that explicitly satisfies these three requirements. The framework adaptively delegates tasks across multiple tiers of expertise: a lightweight base model provides initial answers; a more capable but costlier large model regenerates responses when confidence is low; and, if uncertainty remains high in the model-generated responses, the system abstains to a human expert. An online learning mechanism continually adjusts the deferral and abstention thresholds based on human feedback, improving decision quality over time. Figure 1 provides an overview of this cascaded decision flow with three example questions of varying difficulty.
Contributions.
Our main contributions are threefold: Cascaded LLM Human-AI Decision System: We introduce a multi-tier decision-making system that coordinates LLMs of varying capacity with human experts to balance accuracy, cost, and abstention. Principled Deferral and Abstention Policies: We design confidence- and uncertainty-based decision policies that regulate when to defer to a larger model or abstain to humans, guided by Bayesian calibration for reliable verification. Online Learning for Adaptive Decision-Making: We propose an online optimisation scheme that refines the deferral and abstention thresholds using human feedback, enabling continual adaptation to task complexity.
@@
Figure 1: Cascaded LLM Human-AI Decision-Making Framework Examples. Given a decision-making problem, the system (1) generates an initial response with a base model, (2) verifies correctness probability, (2.5) defers to a larger model if needed, (3) assesses response uncertainty, and (3.5) abstains to a human expert if necessary. If feedback is available, deferral and abstention modules are adjusted over time. For this system to work efficiently, the modules should uphold three desiderata: D1 the deferral policy regenerates responses only when necessary, D2 the abstention policy escalates to humans only when uncertainty is high, D3 the system continuously improves with feedback.
@@ 2 Related Work
Multi-LLM Answer Generation. Several studies have explored collaborative frameworks that leverage multiple LLMs of varying capacities to enhance both performance and cost-efficiency beyond the capabilities of a single model (Chen et al., 2023; Ding et al., 2024; Aggarwal et al., 2024). Chen et al. (2023) proposed cost-effective strategies such as prompt structuring, model approximation, and cascaded LLM frameworks. Similarly, Ding et al. (2024) introduced an intelligent routing mechanism that dynamically assigns prompts to the most appropriate model. Aggarwal et al. (2024) developed a black-box LLM framework for cost-efficient response generation, formalised as a Partially Observable Markov Decision Process (POMDP), requiring minimal training data. Zhu et al. (2023a) proposed a multiplexer-based approach that balances queries between a small and a large LLM, employing a trained BERT classifier to determine when the smaller model suffices. Šakota et al. (2024) introduced a meta-model-driven selection framework that requires pre-training for optimal query distribution. In a parallel line of research, speculative decoding (Leviathan et al., 2023), employs a lightweight model to generate multiple tokens, which a larger model subsequently verifies.
@@
In contrast to prior research, we propose a multi-tier framework for human-AI collaboration. Rather than relying solely on automation, our approach integrates human intervention when model uncertainty is too high, addressing a gap in previous multi-tier frameworks. Compared to speculative decoding research, our work prioritises the factual correctness of complete responses rather than token-wise distributions, enabling more robust decision-making rather than just fluent text generation. Zellinger et al. (2025) conducts a concurrent line of research that is closest to our work on cascaded LLMs, as well as in their previous works (Zellinger and Thomson, 2024, 2025). They focus on probabilistic modelling of cascading LLMs and their deferral and abstention mechanisms.
@@
LLM Answer Verification and Uncertainty Quantification. Ensuring the reliability of LLM-generated responses requires adequate verification and uncertainty quantification mechanisms. Several studies have explored self-verification strategies (Weng et al., 2023; Jiang et al., 2024; Pan et al., 2024), often leveraging the LLM’s internal knowledge (Dhuliawala et al., 2023). Alternative approaches employ external knowledge sources for verification (Pan et al., 2024; Gao et al., 2023; Peng et al., 2023). Aggarwal et al. (2024) introduced verification techniques based on available contextual information, predominantly involving multiple LLM queries to validate response accuracy. Another research direction quantifies factual correctness uncertainty (Mahaut et al., 2024). Kadavath et al. (2022) conducted a detailed analysis of how LLMs express uncertainty through surrogate token probabilities, demonstrating their effectiveness in calibration. Azaria and Mitchell (2023) explored internal LLM states, training classifiers to quantify uncertainty, while methods such as semantic uncertainty estimation (Kuhn et al., 2023) enhance robustness by analysing variations in semantically equivalent token sequences.
@@
LLMs in Online Learning. Traditional LLM research predominantly evaluates language models on static datasets. However, our work aligns with online learning paradigms, wherein policies are continuously refined in response to streaming data (Cortes et al., 2018; Ye et al., 2024). Our methodology is inspired by Jarrett et al. (2022), who introduced an online decision mediation framework mediating between suboptimal human decisions and an expert oracle. A similar research with the online learning approach is conducted by Zhu et al. (2023a), which extended their multiplexer mechanism to an online setting.
@@ 3.1 Cascaded Decision System
[…]
@@
Figure 2: Decision flow of the two-tiered cascaded LLM system. The base model first evaluates each query. Confident, low-uncertainty responses are accepted; uncertain ones are passed to the large model or, if still uncertain, deferred to a human expert. Online feedback progressively improves these policies.
@@ 3.2 Cost Calculation
To estimate the computational cost of response generation, we define a cost function that scales linearly with model size and token counts. Let $s$ denote the model size (in billions of parameters, e.g., Llama3.1-8B $\Rightarrow s=8$), $t_{\text{in}}$ and $t_{\text{out}}$ the numbers of input and output tokens, and $\rho>0$ the output-to-input token cost ratio accounting for the higher cost of generation. The total cost is given by:
@@
This provides a simple yet effective way to compare models of different sizes under a unified cost metric, independent of infrastructure-specific pricing. Additional cost components can be incorporated as needed.
@@ 4.1 Calibrated Confidence and Uncertainty Estimation
For each of the above methods, the evaluating model $\mathcal{M}_{i}$ can either be the same model that generated the original response, or a larger model in the cascade, if available. While self-evaluation is cheap and self-contained, verifying a small model’s output using a larger model is still substantially cheaper than generating a new response from scratch—particularly when generation involves long-form reasoning, as per Assumption A2. Additionally, larger models tend to be better calibrated and may yield more reliable verification, improving downstream deferral and abstention decisions (Zhu et al., 2023b; Chhikara, 2025).
@@
To ensure that the extracted confidence scores are comparable across models and consistent with empirical correctness, we fit a Bayesian logistic regression model on a small calibration set of 100 samples. This is a Bayesian version of Platt scaling (Platt, 2000), and we assume a Normal distribution as prior. We follow Zellinger and Thomson (2024)’s approach and apply a non-linear transformation on the raw confidence score before inputting it into the Bayesian model, to spread out the clusters of overconfident probabilities.
@@ 4.2 Online Improvement
$\mathbb{P}(\text{abstention})\;=\;p_{\text{abst1}}\;+{p_{\text{abst2}}}.$ (12)
@@
$\mathbb{E}[\text{Correct}]\;=\;\mathbb{E}\!\bigl[m_{\text{pred1}}\cdot\Phi_{\text{base}}\bigr]\;+\;\mathbb{E}\!\bigl[m_{\text{pred2}}\cdot\Phi_{\text{large}}\bigr].$ (13)
@@
$\mathbb{E}[\text{Cost}]\;=\;c_{1}\;+\;\mathbb{E}[m_{\text{defer1}}]\cdot c_{2}.$ (14)
@@ 5 Experiments
To evaluate the generalisability of our framework across domains, we use five question-answering datasets: (1) ARC2-Easy and (2) ARC2-Challenge (Clark et al., 2018), which are part of the AI2 Reasoning Challenge and require reasoning over grade-school science; (3) Massive Multitask Language Understanding (MMLU) benchmark (Hendrycks et al., 2021), which covers 57 subjects ranging from complex STEM to international law, nutrition, and religion; and two medical QA benchmarks: (4) MedQA (Jin et al., 2020), consisting of US medical board exam questions, and (5) MedMCQA (Pal et al., 2022), comprising entrance exam questions from the Indian medical school curriculum. All datasets are in multiple-choice format, with ground-truth answers satisfying Assumption A3. Chain-of-Thought reasoning is employed to generate answers. The cost proportion between input and output tokens is set to $\rho=5$, consistent with Anthropic’s current pricing to date (Anthropic, 2025). Details on generation and verification prompts can be found in Appendix B.2.
@@ 5.1 Cost-Benefit Analysis of Verification Methods
To assess both cost-efficiency and accuracy, we compare the calibrated base model confidence $\Phi_{\text{base}}$ against two baselines: (1) using only the base model (Qwen-2.5-1.5B) and (2) using only the large model (Qwen-2.5-7B). In Figure 3, we visualise accuracy versus cost per sample across the datasets. We use a threshold-agnostic strategy where deferral to the large model is performed with probability $\Phi_{\text{base}}(x)$. We evaluate four methods: Self-Verification (SV, n=1n{=}1), Surrogate Token Probability (STP, n=1n{=}1), Consistent Self-Verification (SV, n=5n{=}5), and Monte Carlo STP (MC-STP, n=5n{=}5). For the latter two, we perform five regenerations or stochastic passes. Each experiment is conducted once using $\mathcal{M}_{\text{base}}$ as the verifying LLM, and once using $\mathcal{M}_{\text{large}}$.
@@
As shown in Figure 3 (and Figures 7, 10, and 13 in Appendix C), using a larger model for verification generally yields a better cost-benefit profile, particularly on simpler datasets (ARC2-Easy, ARC2-Challenge, MMLU). In contrast, base-model verification provides only marginal gains. On the more complex medical datasets (MedQA and MedMCQA), all methods struggle. STP (n=1n{=}1) is the most effective the ARC2-Easy and ARC2-Challenge dataset.
@@
ARC2 Easy ARC2 Challenge MMLU MedQA MedMCQA
Base SV (n=1n{=}1) -2.4 $\pm$ 32.6 -11.6 $\pm$ 20.5 -18.7 $\pm$ 4.9 -27.3 $\pm$ 14.3 -24.8 $\pm$ 9.5
SV (n=5n{=}5) 10.2 $\pm$ 29.6 -26.2 $\pm$ 19.1 -19.4 $\pm$ 4.9 -25.4 $\pm$ 13.4 -19.5 $\pm$ 9.4
STP (n=1n{=}1) -16.9 $\pm$ 33.2 -19.3 $\pm$ 20.9 -20.4 $\pm$ 4.5 -23.9 $\pm$ 14.7 -22.9 $\pm$ 9.7
MC-STP (n=5n{=}5) -37.4 $\pm$ 28.8 -16.0 $\pm$ 21.0 -24.6 $\pm$ 4.2 -27.3 $\pm$ 13.7 -22.2 $\pm$ 9.4
Large SV (n=1n{=}1) 258.4 $\pm$ 39.1 70.5 $\pm$ 26.0 -7.8 $\pm$ 4.7 -6.1 $\pm$ 15.8 -11.6 $\pm$ 10.0
SV (n=5n{=}5) 188.5 $\pm$ 31.5 51.0 $\pm$ 23.7 -24.1 $\pm$ 3.9 -11.9 $\pm$ 14.7 -11.0 $\pm$ 9.8
STP (n=1n{=}1) 242.7 $\pm$ 36.3 89.3 $\pm$ 26.2 2.5 $\pm$ 5.2 1.7 $\pm$ 16.3 -10.0 $\pm$ 10.1
MC-STP (n=5n{=}5) 171.4 $\pm$ 36.4 39.2 $\pm$ 22.0 -22.4 $\pm$ 4.0 -16.8 $\pm$ 14.5 -19.7 $\pm$ 9.3
[…]